OpenStudy (kaydra14):

The volume of a rectangular prism can be computed by using the formula V = lwh. Solve the formula for h.

OpenStudy (drummergirl3):

ake everything except r to the other side 2/(2 pie h) = r so, r = s/(2 pie h) 2. a(c - b) = d ok, solve for c, gotta move everything to the other side again. let's start with a (c - b) = d/a now you can move b c = d/a + b taking LCM as a in the right side, c = (d + ab)/a 3. P = 2l + 2w solve for w again, take'em to the other side P - 2l = 2w (P - 2l)/2 = w 4. V = lwh solve for l i think you can do it yourself by now? 5. V = lwh V = 1344 cm^3 w = 8 cm h = 12 cm use the formula you obtained in number 4. and the values given here to calculate l

OpenStudy (kaydra14):

i dont get it